OpenStudy (anonymous):

Wanda is using a punch recipe that calls for 2 liters of club soda and three 750 milliliter bottles of sparkling apple cider. How much of these two ingredients does the recipe call for?

OpenStudy (anonymous):

A 4 liters 2 deciliters 5 centiliters B 5 liters 2 centiliters 5 milliliters C 4 liters 2 centiliters 5 milliliters

OpenStudy (anonymous):

what took u so long

jimthompson5910 (jim_thompson5910):

3*750 mL = 2,250 mL 2 L = 2000 mL ------------------------------------------------------- 2 L + 3*750 mL = 2000 mL + 2250 mL 2 L + 3*750 mL = 4250 mL 2 L + 3*750 mL = 4000 mL + 200 mL + 50 mL 2 L + 3*750 mL = 4 L + 2 cL + 50 dL 2 L + 3*750 mL = 4 Liters, 2 centiliters + 50 decilters and there's no need to be rude...

jimthompson5910 (jim_thompson5910):

oops meant to say 2 L + 3*750 mL = 4 Liters, 2 centiliters, 5 decilters